OpenStudy (anonymous):

will give medal for answer. the perimeter of a rectangle is twice the sum of its length and its width. the perimeter is 16 meters, and its length is 2 meters more then twice its width. What is its length?

OpenStudy (anonymous):

@EmogirlAtEmoooocow

OpenStudy (anonymous):

@Gabby.Rokks

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

Give me some time to figure it out....

OpenStudy (anonymous):

okay @Gabby.Rokks

OpenStudy (anonymous):

i believe its 8

OpenStudy (emogirlatemoooocow):

P = 16 = 2 (l+w) l = 2 + 2w solve :)

OpenStudy (emogirlatemoooocow):

then this 16 = 2 (l+w), so 8 = l +w , so w= 8-l plug this in l = 2 + 2w

OpenStudy (emogirlatemoooocow):

u will get you answer

OpenStudy (anonymous):

thank you @EmogirlAtEmoooocow

OpenStudy (emogirlatemoooocow):

welcome